OpenStudy (sissyedgar):

A silver piece of art in the shape of a regular heptagon has an apothem length of 27 mm and a side length of 26 mm. What is its area?

Directrix (directrix):

The area of a regular polygon is given by the formula A = (1/2)*a*p where a is the apothem of the polygon and p is the perimeter of the polygon A regular heptagon has 7 sides of the same length, 26

OpenStudy (sissyedgar):

ty!

Directrix (directrix):

A = = (1/2)*a*p A = (1/2) * 27 * 7 *26 Solve for A. Post what you get, okay?

Directrix (directrix):

@sissyedgar

OpenStudy (sissyedgar):

im here....

Directrix (directrix):

Multiply 27 * 7 *26. Divide the product by 2. Post what you get.

Directrix (directrix):

Copy and paste this into the search box for Google. The Google calculator will pop up with the answe. (1/2) * 27 * 7 *26=

OpenStudy (sissyedgar):

2,457? {srry i froze}

Directrix (directrix):

Correct. 2457 square mm

OpenStudy (sissyedgar):

oo, ok! thx!

Directrix (directrix):

You are welcome.

OpenStudy (sissyedgar):

another one plz?

Directrix (directrix):

Close this thread and post the new question in a new thread.